# Impacts of aspect ratio on task accuracy in parallel coordinates

### v1

@@ Abstract
Parallel coordinates plots (PCPs) are a widely used visualization method, particularly for exploratory analysis. Previous studies show that PCPs perform much more poorly for estimating positive correlation than for estimating negative correlation [1], but it is not clear if this is affected by the aspect ratio (AR) of the axes pairs. In this paper, we present the results from an evaluation of the effect of the aspect ratio of axes in static (non-interactive) PCPs for two tasks: a) linear correlation estimation and b) value tracing. For both tasks we find strong evidence that AR influences accuracy, including ARs greater than 1:1 being much more performant for estimation of positive correlations. We provide a set of recommendations for visualization designers using PCPs for correlation or value-tracing tasks, based on the data characteristics and expected use cases.
Index Terms:
visualization, perception, parallel coordinates, aspect ratio
I Introduction
PCPs are valuable for multivariate data as they scale effectively with number of attributes displayed [2, 3], with increasing horizontal display space occupied as number of attributes increases (assuming vertical axes). PCPs are used in a a diverse range of fields and applications particularly in multi-dimensional exploratory analysis are wide-ranging [4]. These tasks are comprised of a number of fundamental parts however, notably trends, value-tracing (or outlier detection), correlation estimation and cluster identification [5, 2, 6, 7, 4]. While there are numerous studies evaluating the performance of PCPs for these fundamental tasks (e.g. [1, 8]), little consideration is given to real-world use with differing numbers of axes and layout choices. In practice, the separation distance between PCP axes are frequently set simply by available horizontal screen space and the number of axes required for display [2], or by using a combination of metrics [7]. There are few recommendations for minimum or maximum separation distance, or for aspect ratio (AR, the ratio of the axis separation distance to axis length).
@@ I Introduction
For example, Gehlenborg & Wong [10] note that AR affects the perception of line crossings and suggest that axis height and separation distance should be adjusted such that the absolute value of all angles is close to 45 degrees.
In existing evaluations of PCPs, an AR of 1:1 (1) or 1:2 (0.5) is frequently used. For example, Li et al. [8] compared the perception of correlation levels between scatter plots and PCPs, with both PCPs and scatter plots having a 1:1 ratio.
To the best of our knowledge, there is no existing quantitative evaluation of the effect of the AR on user tasks. It is plausible that differing ARs may affect task success, especially given the differing patterns formed by positive and negative correlations and the associated performance difference [1, 8].
@@
  • Evaluation to determine whether AR has an effect on task success for value tracing and correlation estimation tasks in static (non-interactive) PCPs. Evidence shows that AR does impact on task success, with effect size dependent on both number of samples and AR.
  • Provides evidence that correlation estimation is affected asymmetrically - that is, dependent on the sign of the correlation, with positive correlations being much more accurately estimated at AR $>$1.
  • Recommendations for designers to set AR of axes dependent on expected primary tasks.
The remainder of the paper is structured as follows: Section II presents related work on aspect ratio in visualizations. Hypotheses are described in Section III. The methodology used in our study is described in Section IV, followed by presentation of results in Section V. We discuss the results and limitations of the study in more detail in section VI, followed by future directions in Section VIII.
II Prior work
While there exists extensive work that examines the effect of plot height or AR, it largely focuses on single channels or line plots [11, 12, 13, 14]. Cleveland and McGill introduce the principle of ”banking to 45 degrees” [12] - ensuring that the average slope of a line in a line chart is 45 degrees - which has been frequently used as a guideline in optimization of AR. Based on this, a number of methods to optimise the AR of plots exist [15, 16]. However, this may not always be optimal or minimise error [17, 18].
Figure 1: Correlation in a PCP: (a) patterns formed with increasing correlation strength; (b) patterns in a PCP by AR (central axes - top: positive correlation; bottom: negative correlation)
@@ II Prior work
Decrease in relative line length with increasing AR may impact on slope estimation - Kennedy et al. [22] find that line length has an effect on angle estimation, with most accuracy when line pair length is approximately equal.
@@
Evidence suggests that AR has effects on measures other than slope. For example, Heer and Bostock [14] find that AR affects area estimation.
III Hypotheses
@@
The aim of the research presented in this paper is to investigate whether there is a difference in task success at different axis ARs in PCP. As such, we do not propose perceptual mechanisms for such a difference. Nevertheless, there are some indications from studies of the effect of AR on other visualization methods that a difference may exist for the tasks we select (correlation estimation and value tracing).
@@
Tasks were selected to represent common use-cases for a PCP in a exploratory visualization context [5]. Three tasks are commonly used in evaluations of a PCP - correlation estimation, value tracing and cluster identification [23]. Of these, we examine correlation estimation and value tracing as they have been extensively studied in PCPs previously (with single ARs), providing a strongly validated experimental methodology, and are likely to be substantially affected by AR. Note that we vary AR by keeping height of the PCP consistent, which is appropriate for real-world use cases where a PCP is stretched to fit the full width of the screen regardless of the number of axes. Varying the height of axes is also a potential use-case that may affect perception of patterns, but in order to limit the number of exposure variables it was considered beyond the scope of this initial investigation.
III-A Correlation estimation
@@
In a PCP, negative correlations form an ‘X’ shape, while lines have fewer crossings and become closer to parallel (assuming identical axis scales, i.e. a homogeneous graph) with increasing positive correlation (fig. 1a) [24]. It is likely that distorting the plot horizontally will impact on the perception of these patterns, and, critically, the patterns may be affected differently. There is a negative relationship between AR and between-line angle - as horizontal distance is increased (while maintaining vertical distance), the angle between lines decreases. As such, it is likely that any component of the perceptual mechanism that uses angle estimation will be affected. Acute angles are often overestimated [25] and the tilt effect [26] is likely to apply with estimation of angles in context in a PCP, with bias strongest at around 20-30 degrees. Indeed, Akgöz et al. [27] find that angles between 0 and 10 deg have a strong attraction. Ware et al. [28] suggest that acute angles are more likely to cause visual confusion than angles approaching 90 degrees, based on [29]. Line crossings will likely become less salient at small between-line angles, as is the case at high ARs. For ARs $>$ 1, strongly negative correlations may be less easily perceived (potentially mitigated by the clarity of the pattern). Fig. 1b shows examples of high correlation values at different ARs. Hypothesis 1:
  • H1.1 AR has an effect on accurate estimation of correlation in a PCP.
@@
  • H1.2 AR interacts with correlation to affect accuracy of estimated correlation.
@@
  • H1.3 estimated correlation will more closely match the real value in PCPs with a high AR ($>$1) than in ARs of 1 or lower.
III-B Tracing values
@@
  • H2.1 AR has an effect on successful tracing of values across multiple axes.
  • H2.2 as AR decreases below 1 or increases above 1 successful value tracing will become less likely.
IV Methods
Two tasks were set to evaluate: a) whether AR has an effect on task accuracy and b) the relationship between AR, dataset size and task accuracy.
IV-A Procedures
@@
Participants were asked to identify the correlation between a pair of adjacent PCP axes, by responding to whether the correlation was strongly ($+$/—), moderately (++/–), or weakly (+/-) positive/negative, or if it was zero (0).
The experimental setup was based on the study of correlation perception in PCPs and scatter plots by Li et al. [8]. We adopted an identical approach for selecting correlation strengths, using equidistant levels on the theoretically optimal Fisher-z scale. Datasets of 4 variables with a random normal distribution were created for each sample size (40 and 160 samples) for each correlation between the question variables (+/- 0.905 z = +/- 1.5, +/- 0.762 z = +/- 1, +/- 0.462 z = +/- 0.5, 0 z = 0). Static images of the 4-axis PCP were generated for each size/correlation set and these were then scaled for each AR. ARs were chosen to provide a realistic and readable plot on a single screen with 4 variables with a central AR of 1:1, using a log scale to ensure coverage at the ends of the distribution while closely simulating real values around the 1:1 ratio. Five ARs at 0.25:1, 0.5:1, 1:1, 2:1 and 4:1 were then validated with a pilot study of 10 participants, which indicated that responses had appropriate range with minimal clustering at the extremes. To ensure data quality, task length was reduced by running the experiment in two parts, the first with ARs of 0.25:1, 1:1 and 4:1 and the second with ARs of 0.5:1, 1:1 and 2:1. Questions were randomised to mitigate learning effects where participants recognised the pattern in a plot scaled up or down. This was further mitigated by including the question order as a random effect in the analysis.
@@
Datasets of 4 variables with a random uniform distribution (range 0-50) were created for each sample size (10, 20, 40, 160). Following the procedure set out in [30], to minimise confusion the initial and final values were specifically chosen to ensure no other sample values were closer than 3. Following the procedure in Task A, ARs of 0.25:1, 0.5:1, 1:1, 1:2 and 1:4 were selected to characterise the dynamics of changing response with AR, then validated with a pilot of 10 participants which indicated reasonable variance of responses within AR levels. The pilot indicated that the dataset with 160 samples produced very few correct responses, probably due to visual clutter making the task unfeasible at the required screen resolutions. As such, the 160 sample datasets were removed from the final task.
For each sample size, 3 datasets were generated, creating 9 datasets in total. For each of these datasets, static images of PCPs were created for each AR, leading to a total of 45 plots/questions. A scale was printed on the final axis (the query axis), with a tick at intervals of 5.
Following a pilot study that indicated different datasets could create substantial variability in accuracy, we used the same dataset for each AR, generating 5 plots per dataset. To reduce the risk or a learning effect the answer scale was shifted by AR x 40 and question order was randomised on a per-participant basis.
Figure 2: Example task screens (a) Task A: Correlation estimation. (b) Task B: Valute tracing.
@@
The tasks were implemented using the Gorilla online framework (gorilla.sc). Questions used static images of PCPs, with participants either selecting a button for correlation strength (Task A, fig. 2, top) or inputting the estimated value in a text entry field (Task B, fig. 2, bottom). Question order was randomised on a per-participant basis. For each task, participants were first given guidance on the task (for example correlation patterns in a PCP) and then completed a set of training questions. Training questions asked participants to estimate the correct answer and then click a button to reveal the correct answer. Training questions were formatted identically to the live questions, but with some additional guidance to ensure participant comprehension (Task A). Participants were free to spend as long as they wished completing the training questions. Once training questions were complete, the participants were asked to begin the live questions. Participants were asked 42 questions for Task A and 45 questions for Task B.
@@
Three ‘attention check’ questions were asked interspersed evenly between the trial questions. Where participants incorrectly answered one or more of these questions, all responses from that participant were removed. Where individual responses had a response time of less than 1000ms, the response was removed from further analysis. For Task A, this resulted in the removal of 102 out of 8400 responses (involving 35 participants), resulting in 8298 responses for analysis. For Task B, no responses were removed, resulting in 2160 responses for analysis.
@@
Participants were recruited using the online platform Prolific (prolific.co) and were screened to ensure normal or corrected-normal vision. Participants were asked for consent then asked to rate their experience with visualization (scale 0-5) and statistics (scale 0-5). Demographic data and responses for these questions are detailed in tables I and II. Participants were paid £1.60 for participation in Task A and £3 for participation in Task B. To prevent learning effects, participants that completed the first part of Task A (with ARs 0.25, 1 and 4) were not eligible to complete the second part. For Task A, 205 participants (105 for part 1 and 100 for part 2) were recruited with 200 remaining after QC (103 for part 1 and 97 for part 2); for Task B 49 participants were recruited with 48 remaining after QC.
Participants were asked their age range for both tasks: for Task A the majority of participants were under 40 (86%); for Task B 87.6% were under 40. For further details, see tables I and II. Participants were also asked a series of questions asking for details of background knowledge of statistics and visualization. For Task A, participants were asked to a rate on a Likert scale (1-5) a) their knowledge and experience of visualization methods (median response 3, 75.5% 3 or lower) and b) their knowledge and experience of statistical methods and analysis (median response 3, 74% 3 or lower). See table I for details. For Task B, participants were asked for their years of experience in a) using and/or knowledge of data visualization methods (median 0-2 years, 77.1% less than 2 years) and b) using and/or knowledge of statistical methods and analysis (median 0-2 years, 70.8% less than 2 years). Participants were also asked for their experience of using Parallel Coordinates, with 72.9% having no experience. See table II for details.
TABLE I: Task A - Participant information
Response % participants
Age range 18-25 $35.5$
25-29 $24.5$
30-39 $26.0$
40-49 $9.0$
50-59 $3.0$
60-69 $1.0$
70-79 $1.0$
80+ $0.0$
Response % participants
Knowledge and experience of data visualization methods 1 $11.0$
2 $24.0$
3 $40.5$
4 $19.0$
5 $5.5$
Knowledge of statistical methods and analysis 1 $13.5$
2 $22.0$
3 $38.5$
4 $20.0$
5 $6.0$
TABLE II: Task B - Participant information
Response % participants
Age range 18-25 $29.2$
25-29 $41.7$
30-39 $16.7$
40-49 $12.5$
50-59 $0.0$
60-69 $0.0$
70-79 $0.0$
80+ $0.0$
Response % participants
Years of experience of using and/or knowledge of data visualization methods None $39.6$
0-2 years $37.5$
2-5 years $18.8$
5-10 years $4.2$
10-20 years $0.0$
20+ years $0.0$
Years of experience of using and/or knowledge of statistical methods and analysis None $35.4$
0-2 years $35.4$
2-5 years $16.7$
5-10 years $12.5$
10-20 years $0.0$
20+ years $0.0$
Response % participants
Experience of using Parallel Coordinates None $72.9$
Used occasionally $27.1$
Used frequently $0.0$
IV-B Analysis methodology

### v2

@@ Abstract
Parallel coordinates plots (PCPs) are a widely used visualization method, particularly for exploratory analysis. Among the strengths of PCPs are their ability to support diverse tasks in multivariate data analysis, such as correlation estimation, value tracing and cluster identification. Previous studies show that PCPs perform much more poorly for estimating positive correlation than for estimating negative correlation [1], but it is not clear if this is affected by the aspect ratio (AR) of the axes pairs. In this paper, we present the results from an evaluation of the effect of the aspect ratio of axes in static (non-interactive) PCPs for two tasks: a) linear correlation estimation and b) value tracing. For both tasks we find strong evidence that aspect ratio influences accuracy, including aspect ratios greater than 1:1 being much more performant for estimation of positive correlations. We provide a set of recommendations for visualization designers using PCPs for correlation or value-tracing tasks, based on the data characteristics and expected use cases.
keywords:
parallel coordinates, aspect ratio, correlation
1 Introduction
PCPs are valuable for multivariate data as they scale effectively with number of attributes displayed [2, 3], with increasing horizontal display space occupied as number of attributes increases (assuming vertical axes). PCPs are used in a diverse range of fields and applications particularly in multi-dimensional exploratory analysis [4]. These tasks are comprised of a number of fundamental parts, notably trends, value-tracing (or outlier detection), correlation estimation and cluster identification [5, 2, 6, 7, 4]. While there are numerous studies evaluating the performance of PCPs for these fundamental tasks (e.g. [1, 8]), little consideration is given to real-world use with differing numbers of axes and layout choices. In practice, the separation distance between PCP axes are frequently set simply by available horizontal screen space and the number of axes required for display [2], or by using a combination of metrics [7]. There are few recommendations for minimum or maximum separation distance, or for aspect ratio (AR, the ratio of the horizontal axis separation distance to axis height).
@@ 1 Introduction
For example, Gehlenborg & Wong [10] note that aspect ratio affects the perception of line crossings and suggest that axis height and separation distance should be adjusted such that the absolute value of all angles is close to 45 degrees.
In existing evaluations of PCPs, an aspect ratio of 1:1 (1) or 1:2 (0.5) is frequently used. For example, Li et al. [8] compared the perception of correlation levels between scatter plots and PCPs, with both PCPs and scatter plots having a 1:1 ratio.
To the best of our knowledge, there is no existing robust quantitative evaluation of the effect of aspect ratio on user tasks. It is plausible that differing aspect ratios may affect task success, especially given the differing patterns formed by positive and negative correlations and the associated performance difference [1, 8].
In parallel to our work, Meka et al. [11] conducted a study to explore simple negative/positive correlation perception in PCPs under varying aspect ratios and correlation coefficients. This study does not appear to find significant differences in accuracy of responses with the exception of 1:1 and 9:16. The response measure simply evaluated whether the participants correctly identified negative or positive correlations and as such inferences are very limited, with no consideration of the under- or over-estimation of the strength of a correlation.
@@
  • Evaluation to determine whether aspect ratio has an effect on task success for value tracing and correlation estimation tasks in static (non-interactive) PCPs. Evidence shows that aspect ratio does impact on task success, with effect size dependent on both number of samples and aspect ratio.
  • Provides evidence that correlation estimation is affected asymmetrically - that is, dependent on the sign of the correlation, with positive correlations being much more accurately estimated at aspect ratio >>1.
  • Recommendations for designers to set aspect ratio of axes dependent on expected primary tasks.
The remainder of the paper is structured as follows: Section 2 presents related work on aspect ratio in visualizations. Hypotheses are described in Section 3. The methodology used in our study is described in Section 4, followed by presentation of results in Section 5. We discuss the results and limitations of the study in more detail in section 6, followed by future directions in Section 8.
2 Prior work
While there exists extensive work that examines the effect of plot height or aspect ratio, it largely focuses on single channels or line plots [12, 13, 14, 15]. Cleveland and McGill introduce the principle of ”banking to 45 degrees” [13] - ensuring that the average slope of a line in a line chart is 45 degrees - which has been frequently used as a guideline in optimization of aspect ratio. Based on this, a number of methods to optimise the aspect ratio of plots exist [16, 17]. However, this may not always be optimal or minimise error [18, 19].
Figure 1: Correlation in a PCP: (a) patterns formed with increasing correlation strength; (b) patterns in a PCP by aspect ratio (central axes - top: positive correlation; bottom: negative correlation)
@@ 2 Prior work
Decrease in relative line length with increasing aspect ratio may impact on slope estimation - Kennedy et al. [23] find that line length has an effect on angle estimation, with most accuracy when line pair length is approximately equal.
@@
Evidence suggests that aspect ratio has effects on measures other than slope. For example, Heer and Bostock [15] find that aspect ratio affects area estimation.
This provides a strong indication that features affected by aspect ratio will impact effectiveness of a visualization, but there is no work that directly evaluates the strength of any effect on a PCP.
3 Hypotheses
@@
The aim of the research presented in this paper is to investigate whether there is a difference in task success at different axis aspect ratios in PCP. Given the specific aim of understanding task success, we do not propose perceptual mechanisms for such a difference. Nevertheless, there are some indications from studies of the effect of aspect ratio on other visualization methods that a difference may exist for the tasks we focus on (correlation estimation and value tracing).
@@
Tasks were selected to represent common use-cases for a PCP in a exploratory visualization context [5]. Three tasks are commonly used in evaluations of a PCP - correlation estimation, value tracing and cluster identification [24]. Of these, we examine correlation estimation and value tracing as they have been extensively studied in PCPs previously (with single aspect ratios), providing a strongly validated experimental methodology, and are likely to be substantially affected by aspect ratio. Note that we vary aspect ratio by keeping the height of the PCP consistent, which is appropriate for real-world use cases where a PCP is stretched to fit the full width of the screen regardless of the number of axes. Varying the height of axes is also a potential use-case that may affect perception of patterns, but in order to limit the number of exposure variables it was considered beyond the scope of this initial investigation.
3.1 Correlation estimation
@@
In a PCP, negative correlations form an ‘X’ shape, while lines have fewer crossings and become closer to parallel (assuming identical axis scales, i.e. a homogeneous graph) with increasing positive correlation (fig. 1a) [25]. It is likely that distorting the plot horizontally will impact on the perception of these patterns, and, critically, the patterns may be affected differently. There is a negative relationship between aspect ratio and between-line angle - as horizontal distance is increased (while maintaining vertical distance), the angle between lines decreases. As such, it is likely that any component of the perceptual mechanism that uses angle estimation will be affected. Acute angles are often overestimated [26] and the tilt effect [27] is likely to apply with estimation of angles in context in a PCP, with bias strongest at around 20-30 degrees. Indeed, Akgöz et al. [28] find that angles between 0 and 10 deg have a strong attraction. Ware et al. [29] suggest that acute angles are more likely to cause visual confusion than angles approaching 90 degrees, based on [30]. Line crossings will likely become less salient at small between-line angles, as is the case at high aspect ratios. For aspect ratios >> 1, strongly negative correlations may be less easily perceived (potentially mitigated by the clarity of the pattern). Fig. 1b shows examples of high correlation values at different aspect ratios. Hypothesis 1:
  • H1.1 Aspect ratio has an effect on accurate estimation of correlation in a PCP.
@@
  • H1.2 Aspect ratio interacts with correlation to affect accuracy of estimated correlation.
@@
  • H1.3 estimated correlation will more closely match the real value in PCPs with a high aspect ratio (>>1) than in aspect ratios of 1 or lower.
3.2 Tracing values
@@
  • H2.1 Aspect ratio has an effect on successful tracing of values across multiple axes.
  • H2.2 as aspect ratio decreases below 1 or increases above 1 successful value tracing will become less likely.
4 Methods
Two tasks were set to evaluate: a) whether aspect ratio has an effect on task accuracy and b) the relationship between aspect ratio, dataset size and task accuracy.
4.1 Procedures
@@
Participants were asked to identify the correlation between a pair of adjacent PCP axes, by responding to whether the correlation was strongly, moderately, or weakly positive/negative, or if it was zero.
The experimental setup was based on the study of correlation perception in PCPs and scatter plots by Li et al. [8]. We adopted an identical approach for selecting correlation strengths, using equidistant levels on the theoretically optimal Fisher-z scale. The Fisher-z transformation transforms the sampling distribution of Pearson’s r into a normal distribution, reducing skew at high values of r. Li et al. [8] find that this theoretically optimal scale is a reasonable reflection of the human perceptual scale for PCPs. We note the relevant Pearson’s r values associated with the selected Fisher-z scores in the parentheses below. Datasets of 4 variables with a random normal distribution were created for each sample size (40 and 160 samples) for each correlation between the question variables ($r$ = +/- 0.905 z = +/- 1.5; $r$ = +/- 0.762 z = +/- 1; $r$ = +/- 0.462 z = +/- 0.5; $r$ = 0 z = 0). Static images of the 4-axis PCP were generated for each size/correlation set and these were then scaled for each aspect ratio. Buttons were displayed below the plot, with a magnitude indicator for strong (+++/—), moderate (++/–), weak (+/-) or zero (0) correlation (see fig. 2a). Participants selected a button with the appropriate magnitude and sign of correlation as their response.
Aspect ratios were chosen to provide a realistic and readable plot on a single screen with 4 variables with a central aspect ratio of 1:1, using a log scale to ensure coverage at the ends of the distribution while closely simulating real values around the 1:1 ratio. Five aspect ratios at 0.25:1, 0.5:1, 1:1, 2:1 and 4:1 were then validated with a pilot study of 10 participants, which indicated that responses had appropriate range with minimal clustering at the extremes. To ensure data quality, task length was reduced by running the experiment in two parts, the first with aspect ratios of 0.25:1, 1:1 and 4:1 and the second with aspect ratios of 0.5:1, 1:1 and 2:1. Questions were randomised to mitigate learning effects where participants recognised the pattern in a plot scaled up or down. This was further mitigated by including the question order as a random effect in the analysis.
@@
Datasets of 4 variables with a random uniform distribution (range 0–50) were created for each sample size (10, 20, 40, 160). Following the procedure set out in [31], to minimise confusion the initial and final values were specifically chosen to ensure no other sample values were closer than 3. Following the procedure in Task A, aspect ratios of 0.25:1, 0.5:1, 1:1, 2:1 and 4:1 were selected to characterise the dynamics of changing response with aspect ratio, then validated with a pilot of 10 participants which indicated reasonable variance of responses within aspect ratio levels. The pilot indicated that the dataset with 160 samples produced very few correct responses, probably due to visual clutter making the task unfeasible at the required screen resolutions. As such, the 160 sample datasets were removed from the final task.
For each sample size, 3 datasets were generated, creating 9 datasets in total. For each of these datasets, static images of PCPs were created for each aspect ratio, leading to a total of 45 plots/questions. A scale was printed on the final axis (the query axis), with a tick at intervals of 5.
Following a pilot study that indicated different datasets could create substantial variability in accuracy, we used the same dataset for each aspect ratio, generating 5 plots per dataset. To reduce the risk or a learning effect the answer scale was shifted by aspect ratio x 40 and question order was randomised on a per-participant basis.
Figure 2: Example task screens (a) Task A: Correlation estimation. (b) Task B: Value tracing.
@@
The tasks were implemented using the Gorilla online framework (gorilla.sc). Questions used static images of PCPs, with participants either selecting a button for correlation strength (Task A, fig. 2a) or inputting the estimated value in a text entry field (Task B, fig. 2b). Question order was randomised on a per-participant basis. For each task, participants were first given guidance on the task (for example correlation patterns in a PCP) and then completed a set of training questions. Training questions asked participants to estimate the correct answer and then click a button to reveal the correct answer. Training questions were formatted identically to the live questions, but with some additional guidance to ensure participant comprehension (Task A). Participants were free to spend as long as they wished completing the training questions. Once training questions were complete, the participants were asked to begin the live questions. Participants were asked 42 questions for Task A and 45 questions for Task B.
@@
Three ‘attention check’ questions were asked interspersed evenly between the trial questions. Where participants incorrectly answered one or more of these questions, all responses from that participant were removed. Responses under 850ms were removed based on tests of minimum response times and the distribution of real responses. For Task A, this resulted in the removal of 43 out of 8400 responses (involving 13 participants), resulting in 8298 responses for analysis. For Task B, no responses were removed, resulting in 2160 responses for analysis.
@@
Participants were recruited using the online platform Prolific (prolific.co) and were screened to ensure normal or corrected-normal vision. Participants were asked for consent then asked to rate their experience with visualization (scale 0-5) and statistics (scale 0-5). Demographic data and responses for these questions are detailed in the supplementary material. Participants were paid £1.60 for participation in Task A and £3 for participation in Task B. To prevent learning effects, participants that completed the first part of Task A (with aspect ratios 0.25, 1 and 4) were not eligible to complete the second part. For Task A, 205 participants (105 for part 1 and 100 for part 2) were recruited with 200 remaining after quality control (QC) (103 for part 1 and 97 for part 2); for Task B 49 participants were recruited with 48 remaining after QC.
Participants were asked their age range for both tasks: for Task A the majority of participants were under 40 (86%); for Task B 87.6% were under 40. For further details, see supplementary materials. Participants were also asked a series of questions asking for details of background knowledge of statistics and visualization. For Task A, participants were asked to a rate on a Likert scale (1-5) a) their knowledge and experience of visualization methods (median response 3, 75.5% 3 or lower) and b) their knowledge and experience of statistical methods and analysis (median response 3, 74% 3 or lower).
For Task B, participants were asked for their years of experience in a) using and/or knowledge of data visualization methods (median 0-2 years, 77.1% less than 2 years) and b) using and/or knowledge of statistical methods and analysis (median 0-2 years, 70.8% less than 2 years). Participants were also asked for their experience of using Parallel Coordinates, with 72.9% having no experience.
4.2 Analysis methodology